# Talent Hoarding in Organizations

### v2

@@ 5 Evidence from Workers: Effects of Manager Rotations
This section provides evidence that talent hoarding has meaningful impacts on worker outcomes. I use manager rotations as a source of quasi-random variation in workers’ exposure to talent hoarding. When a manager learns that they will move to a new position on a different team, they no longer have an incentive to hoard workers on their current team. For workers whose manager will soon rotate, this creates a temporary window of time during which they will not be subject to talent hoarding. I test for the impacts of talent hoarding using workers’ internal job applications, which represent high-stakes actions that are critical for workers’ careers within the firm. This quasi-experimental analysis provides evidence that complements the earlier analysis of managers, and offers the advantage of not restricting to a specific measure of hoarding behavior. To support the interpretation of rotation effects as being driven by talent hoarding, Section 5.1 tests predictions from the conceptual framework, while Section 5.2 discusses potentially confounding mechanisms. These empirical exercises support my interpretation based on talent hoarding and motivate the use of rotations to demonstrate the impacts of hoarding on talent allocation in Section 6. Manager rotations represent typical features of managers’ career progression and are not restricted to specific managers or teams. As in many other firms, managers switch positions within the firm to fulfill requirements for future promotions (Cullen and Perez-Truglia, 2023). [^17] I use the term manager rotation in line with previous research to refer to the event, in which a manager leaves their team for a different position within the firm. For instance, in order to become an executive, the firm requires managers to collect experience in different parts of the firm. In general, internal job switches are routine and encouraged by the firm. Consistent with the institutional environment, I empirically document that manager rotations appear orthogonal to pre-existing worker characteristics by comparing workers who are exposed to a manager rotation to those that are not (Panel A of Appendix Table 3). I find that both sets of workers are observationally similar with respect to their demographics and their past performance. Panel B of Table 3 finds a similar balance with respect to manager characteristics, such as demographics, performance, and pre-event team-level outcomes. Appendix Figure A4 uses an event study design to document the absence of pre-trends in key team-level outcomes, such as absenteeism, further corroborating that manager rotations are not restricted to teams facing specific circumstances. My main analysis uses 1,359 manager rotations that result from a worker’s direct supervisor leaving their team to make an internal job transition within the firm. [^18] Appendix Section E presents a robustness test using manager exits, which yields similar patterns. During the four-year study period, 20% of managers rotate at least once. Rotations do not occur on a regular schedule, and workers cannot easily predict when managers will leave the team; Appendix Figure A2 documents large variation in the time that managers spend in a position before rotating. The firm reports that managers typically learn about their new position two to three quarters before they rotate, at which point they inform their teams about the rotation, following a strict policy at the firm that requires them to announce their transitions as soon as possible. [^19] Survey responses, indicating that the news of internal applications spreads quickly by internal gossip, further corroborate the immediacy of these announcements since it is readily apparent when someone has violated this policy. Because managers know they will not be affected by workers leaving the team once they find out about their departure, manager rotations should temporarily reduce hoarding incentives, and thus increase worker applications. I illustrate the dynamic effects of manager rotations using a quarterly event study of applications around the quarter in which a manager rotates. I estimate a specification with worker and quarter fixed effects, binning event time dummy variables at $t=-8$ and $t=4$, and clustering standard errors at the worker and rotation level. Panel A of Figure 3 presents quarterly event study coefficients and demonstrates that manager rotations result in a distinct transitory increase in worker applications. Event time $t=0$ denotes the quarter in which a manager rotates. Application rates increase up to three quarters before the manager rotation takes place, which is when managers start to inform their teams about their departure. However, before $t=-3$, when managers do not know yet about their job rotation, trends in applications are flat. In the quarter of rotation, applications increase by 2.3 percentage points, almost doubling workers’ baseline application rate of 2.7%. As the manager’s replacement settles in, application rates return to baseline levels after one quarter. [^20] Note that applications taking a quarter to return to baseline is consistent with the fact that in a given quarter, only few job openings are available in the internal labor market, implying that not all workers who decide to apply once they find out about the manager rotation will immediately find a suitable job opening. Since manager rotations appear to have the largest impact on worker applications in the quarter of the rotation, the remainder of the analysis focuses on worker behavior in that quarter. I analyze the effect of manager rotations on worker applications in the same quarter by estimating a linear model for workers’ internal application choices using an OLS regression of the following form:
@@
$\text{Applied}_{it}$ and $\text{Rotation}_{it}$ are indicators that worker $i$ in quarter $t$ applies for an internal job opening and experiences a manager rotation, respectively. $X_{it}$ includes a broad set of worker and position controls. [^21] My baseline set of controls, described in Section 3.2, include female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, and number of direct reports. Appendix Table E1 documents that my results are robust to including different controls. $\delta_{t}$ represents quarter fixed effects. Robust standard errors are two-way clustered at the rotation and worker level. Manager rotations have large positive impacts on worker applications in the quarter of rotation. Column 1 of Table 4, Panel A presents OLS estimates for the effect of manager rotations on worker applications, based on Equation 3. When a manager rotates, applications increase by 2.3 percentage points, representing a 78% increase compared to the baseline application rate of 2.9%. The application effect is not a result of managers taking their subordinates with them to the new team or of workers replacing managers in their old position: 97% of applications at the firm are for positions outside of the worker’s current team and not to the manager’s new team. In addition, the observed impact does also not depend on the characteristics of the outgoing or incoming managers (Appendix Table B6). If manager rotations reflect the impact of talent hoarding, these findings imply that managerial talent hoarding deters a meaningful share of applicants from exploring job opportunities outside the team. The following subsection presents evidence in favor of this interpretation. Interpreted in this light, this finding accords with workers’ survey responses discussed in Section 2, in which 41% of workers indicate that they are afraid to apply to internal positions due to talent hoarding.
@@ 5.1 Does Talent Hoarding Drive Rotation Effects?
This section provides evidence that talent hoarding is a key mechanism underlying the impacts of manager rotations. The conceptual framework predicts that different types of workers face different magnitudes of talent hoarding, depending on factors like productivity and manager costs of hoarding. I test these predictions by comparing the effects of rotations on different groups of workers. Worker Quality.—The conceptual framework in Section 2.2 predicts that managers are more likely to hoard high-productivity workers. Therefore, when managers rotate and talent hoarding temporarily subsides, application effects should be larger for the high-productivity workers who experienced more talent hoarding than their low-productivity peers. To test this prediction, I compare the impacts of rotations on applications by worker quality. I construct a measure of worker quality as the predicted values from a regression of applicants’ internal hiring probabilities on applicant characteristics. The predicted value of this regression provides an index of worker quality for all workers, weighting worker characteristics by their importance for hiring prospects within the firm. [^22] This definition of worker quality has the advantage that it reflects the values that the firm places on worker qualifications. Robustness exercises in Appendix Section E find similar patterns when using alternative measures such as educational qualifications or past performance. For instance, while manager rotations increase applications of workers with a graduate degree by 2.8 percentage points, this increase is only 1.3 percentage points for workers without a graduate degree ($p$-value of difference=0.008). I compare workers in the top and bottom quartile of the index. Panel A of Figure 4 shows that high-quality workers experience a 4.4 percentage point increase in applications when a manager rotates, while applications among low-quality workers increase by only 0.8 percentage points. This statistically significant difference ($p$=0.000) is in line with the theoretical predictions. Departure Costs.—Another prediction is that managers hoard talent more when the costs of worker departures are larger. I test this prediction using two measures of departure costs. The first measure is the number of other workers in the team, since a larger team contains more workers that can compensate for a teammate’s departure. Panel B of Figure 4 presents effects of manager rotations for workers in the bottom and top quartiles of team size. By including detailed worker and position characteristics, this analysis compares workers with similar qualifications in similar positions, but in teams of different sizes. Applications increase by 3.5 percentage points among workers in small teams with four or less teammates, but only by 1.7 percentage points in large teams with 10 or more teammates ($p$-value of difference=0.030). Measuring departure costs using the average number of days required to fill a worker’s vacated position yields similar results. Panel C of Figure 4 compares effects for high-cost workers (more than 140 days to fill a vacancy) to low-cost workers (less than 112 days). Applications among workers who are hard to replace increase by 2.3 percentage points, compared to 1.2 percentage points for those who are easier to replace ($p$=0.107). Talent Hoarding Propensity.—The impacts of manager rotations are strongly correlated with direct measures of managers’ propensities to hoard talent. Panel D of Figure 4 compares the impacts of rotations between managers with a high versus low propensity to hoard talent, using my measure of talent hoarding based on the suppression of potential ratings (described in detail in Section 4). Applications increase by 3.4 percentage points when a manager with a high propensity to hoard talent rotates, but only by 1.4 percentage points under managers with a low propensity to hoard talent ($p$-value of difference=0.008). The framework also predicts that hoarding-prone managers have larger impacts on more productive workers. Figure 5 confirms this prediction, indicating that rotation effects of hoarding-prone managers are concentrated among high-quality workers. Risk of Retaliation—In survey results (Section 2.1), workers report fear of manager retaliation as a reason they refrain from applying. If workers are more likely to refrain from applying to positions that carry a greater risk of retaliation, applications to these positions should exhibit particularly large increases around rotations. I test this prediction by comparing workers’ applications to job openings that are closer in proximity to their current position (i.e., within versus outside their current division, functional area, and location). Intuitively, the risk of retaliation is larger if managers are more likely to find out about worker applications. Appendix Table B5 shows that manager rotations have much larger and statistically distinguishable effects on applications in close proximity to workers’ current position with respect to all three dimensions, confirming the prediction. [^23] Manager rotations could lead to an opposite prediction on proximity if they were to operate through the visibility channel instead of retaliation. However, while the suppression of visibility is a common form of talent hoarding, manager rotations do not immediately increase visibility since worker evaluations only occur one to two times a year. External Transitions.—While managers may frequently learn about workers’ applications to internal positions and possibly retaliate or interfere with those applications, this is not the case for applications to positions outside the firm. [^24] In Germany, it is very uncommon that potential employers contact the current firm to inquire about the candidate. Except for a few specific occupations, such as academics, the use of reference letters is also very uncommon, implying that managers have both little ability to learn about or interfere with external applications. Under talent hoarding, manager rotations should thus only impact internal applications and not external applications outside the firm. Because I do not observe external applications, I test this prediction by comparing the effect of manager rotations on internal job transitions within the firm and external job transitions out of the firm. Panel B of Figure 3 shows that manager rotations only increase worker transitions within the firm and not outside of the firm, even though both types of transitions trend identically in quarters prior to the rotation. Columns 1 and 2 of Appendix Table B4 confirm this finding. Manager rotations increase internal transitions by 1.1 percentage points but have negligible effects on external transitions (0.11 percentage points) compared to the identical baseline rate of 0.8% ($p$-value of difference=0.000).
5.2 Testing for Alternative Mechanisms
The analysis in the previous section is consistent with talent hoarding being a key driver of worker responses to manager rotations; however, it remains possible that worker responses could also reflect factors unrelated to talent hoarding. For instance, managers and workers may jump ship because of bad news about the future outlook of the team, or may depart after the completion of a major milestone. Other alternative mechanisms such as worker loyalty, match effects, and role-model effects could also produce the observed worker responses. In this section, I conduct a series of additional robustness checks to assess the potential importance of such alternative mechanisms. Bad news about the future or unpleasant working conditions.— I begin by testing whether a negative shock, such as bad news about the future outlook of the unit or unpleasant working conditions, could underlie the effect of manager rotations on worker applications. Under such a mechanism, one would expect to see effects on other team-level outcomes besides applications, such as team performance and morale. Appendix Figure A4 estimates an event study around manager rotations and rejects economically significant changes in key outcomes, including team-level absenteeism and performance-related bonus pay, suggesting that rotations are not correlated with deteriorating conditions within the team. Results from two placebo tests further contradict patterns that one would expect from such shocks. First, under a correlated shock that makes managers and workers want to leave the team, one would expect increased worker applications even for managers’ unsuccessful attempts to leave the team. However, Panel A of Appendix Figure A5 finds no effect of managers’ unsuccessful rotations on worker applications. Second, because coworkers are more similar in their career opportunities to a worker, one would also expect the correlation between coworker rotations and worker applications in the presence of such a shock to be larger than that between manager rotations and worker applications. Panel B of Appendix Figure A5 estimates the impacts of rotations of the most senior teammate and shows that these events have much smaller impacts on worker applications. In addition, if the documented increase in internal applications is driven by teammates jumping ship, one would expect to see at least some effect on external transitions. However, I do not find that manager rotations yield increases in exit from the firm (Panel B of Figure 3). Moreover, if teammates want to jump ship because of the long-term future outlook of the team, one should expect the increase in applications to persist for several quarters, particularly given that in the internal labor market, it can take several quarters for suitable vacancies to become available. Instead, the observed application responses are transitory. It is also unclear why workers in smaller teams, those that are harder to replace, and those under hoarding-prone managers would be more susceptible to such a shock, as suggested by the heterogeneity results presented in Figure 4.252525Managers’ hoarding propensities are quite stable over time and are unlikely to be driven by a correlated shock. Completion of a milestone or project.—If most teams were to work on fixed project schedules together, the completion of a milestone could be another key channel for correlated shocks. Several institutional features imply that such milestones are not likely to be relevant in this setting. First, manager rotations are not designed to coincide with increases in job vacancies or team dissolution that would indicate systematic rotation patterns due to common milestones in the firm. Second, the vast majority of positions at the firm are not project-based. Even in areas where project-based work is more common, projects are not normally staffed within a team, but often have employees from different teams in one project, meaning that it is not common that an entire team and their manager will finish a milestone at the same time. To empirically test for this channel, I exclude workers in jobs related to project management (where the bulk of project-based work occurs) from my event study analysis. This robustness test yields very similar application patterns (Panel B of Appendix Figure E1), suggesting that my results are not likely to be driven by project cycles. These findings are also consistent with the absence of effects on external transitions. If milestone completion were to drive responses to rotations, one would expect some effect on external transitions. Loyalty.—In principle, manager rotations could induce workers to apply if they feel a strong sense of loyalty toward their manager. While loyalty undoubtedly represents an important part of interpersonal relations in the workplace, I do not find evidence that this mechanism drives the particular increase in applications around manager rotations. Loyalty is typically assumed to compound over time; however, Appendix Figure A3 finds no evidence that rotation effects vary by the length of time a worker was exposed to the rotating manager. Instead, rotations increase applications even for workers who have been exposed to their manager for one quarter or less. While one could have also expected that loyalty is particularly strong if the worker was personally hired by the manager, Appendix Table B4 finds similar effects for workers who were hired by the rotating manager and those who were already in the team. Moreover, one would expect loyalty effects to also result in increases in external transitions. However, Panel B of Figure 3 shows that while internal and external transitions trend identically prior to the rotation, manager rotations do not yield an increase in worker’s external transition out of the firm. Worker-Manager Specific Match Effects.—Workers may have refrained from applying prior to a manager rotation because of match effects, which could occur because the manager provided favorable task assignments or offered better training opportunities. Several pieces of evidence suggest, however, that standard match effects cannot explain the increase in applications around a manager rotation. Similar to loyalty, match effects should increase worker applications for positions outside the team because the value of their default option falls when the manager leaves. However, I find no impact of manager rotations on external transitions. Under match effects, one would also expect rotations to be more impactful for workers who were hired by that manager, relative to workers who were already on the team when that manager arrived. This is because the manager would have had the ability to select new workers based on match quality. Instead, Appendix Table B4 documents that manager rotations have very similar effects on both groups. While one would also expect that match effects compound over time, I find that application effects of manager rotations do not depend on workers’ exposure time to their manager (Appendix Figure A3). Finally, it is difficult to rationalize why match effects would give rise to the observed heterogeneity across different workers, for instance in terms of larger rotation effects for workers under hoarding-prone managers. Evidence discussed in Section 4 suggests that match effects are likely smaller, not larger, under hoarding-prone managers, for example because they deter workers’ participation in trainings. Salience and Role-Model Effects.—Information-based mechanisms could also explain why rotations affect worker applications. For instance, managers who pursue rotations may generate information flows that make career planning more salient. However, since managers’ unsuccessful rotation attempts do not increase applications even though workers likely find out about them, such a channel does not appear to operate through the salience of career planning. Successful manager rotations, on the other hand, could increase applications by providing information about the success likelihood of applications. If this were the case, one would expect that the increase in applications would persist, rather than quickly dissipate, given that few vacancies are available each quarter. Role-model effects are often found to be particularly impactful if role models are similar in attributes to affected individuals (Porter and Serra, 2020, Meier et al., 2020), but homophily does not appear to play an important role in my setting. First, motivated by the previous literature on homophily by gender, I show that the rotation of a manager with the same gender yields a similar and statistically indistinguishable application response relative to the rotation of a manager with the opposite gender (Panel A of Appendix Table B6). Second, if observing others navigate their career is a key underlying factor of rotation effects, such a role-model effect should not be limited to managers. Since coworkers are more similar to workers than their managers with respect to their career opportunities, one would likely expect larger effects for observing coworkers rotate. However, Panel B of Appendix Figure A5 finds much smaller, rather than larger, effects of coworker rotations relative to manager rotations. These findings indicate that any role-model effects that are at play likely operate through channels that differ from those documented in prior research.

### v3

@@ 5 Evidence from Workers: Effects of Manager Rotations
This section provides evidence that talent hoarding has meaningful impacts on worker outcomes. I use manager rotations as a source of quasi-random variation in workers’ exposure to talent hoarding. When a manager learns that they will move to a new position on a different team, they no longer have an incentive to hoard workers on their current team. For workers whose manager will soon rotate, this creates a temporary window of time during which they will not be subject to talent hoarding. I test for the impacts of talent hoarding using workers’ internal job applications, which represent high-stakes actions that are critical for workers’ careers within the firm. This quasi-experimental analysis provides evidence that complements the earlier analysis of managers. While the manager-level analysis isolates specific managerial actions—such as suppressing worker visibility—it captures only one particular hoarding action, namely, the suppression of worker visibility outside of the team. In contrast, the worker-level analysis based on manager rotations encompasses a broader set of hoarding actions, including less easily-measured actions such as discouraging applications or exerting implicit pressure. Section 5.1 introduces the rotation design. To support the interpretation of rotation effects as being driven by talent hoarding, Section 5.2 tests predictions from the conceptual framework.
5.1 Rotation Design
Manager rotations represent typical features of managers’ career progression and are not restricted to specific managers or teams. As in many other firms, managers switch positions within the firm to fulfill requirements for future promotions (Cullen and Perez-Truglia, 2023). [^18] I use the term manager rotation in line with previous research to refer to the event, in which a manager leaves their team for a different position within the firm. For instance, in order to become an executive, the firm requires managers to collect experience in different parts of the firm. In general, internal job switches are routine and encouraged by the firm. Consistent with the institutional environment, I empirically document that manager rotations appear orthogonal to pre-existing worker characteristics by comparing workers who are exposed to a manager rotation to those that are not (Panel A of Appendix Table 4). I find that both sets of workers are observationally similar with respect to their demographics and their past performance. Panel B of Table 4 finds a similar balance with respect to manager characteristics, such as demographics, performance, and pre-event team-level outcomes. Appendix Figure A2 uses an event study design to document the absence of pre-trends in key team-level outcomes, such as absenteeism, further corroborating that manager rotations are not restricted to teams facing specific circumstances.
My main analysis uses 1,359 manager rotations that result from a worker’s direct supervisor leaving their team to make an internal job transition within the firm. [^19] Appendix Section E presents a robustness test using manager exits, which yields similar patterns. During the four-year study period, 20% of managers rotate at least once. Rotations do not occur on a regular schedule, and workers cannot easily predict when managers will leave the team; Appendix Figure A1 documents large variation in the time that managers spend in a position before rotating. The firm reports that managers typically learn about their new position two to three quarters before they rotate, at which point they inform their teams about the rotation, following a strict policy at the firm that requires them to announce their transitions as soon as possible. [^20] Survey responses, indicating that the news of internal applications spreads quickly by internal gossip, further corroborate the immediacy of these announcements since it is readily apparent when someone has violated this policy.
Because managers know they will not be affected by workers leaving the team once they find out about their departure, manager rotations should temporarily reduce hoarding incentives, and thus increase worker applications. I illustrate the dynamic effects of manager rotations using a quarterly event study of applications around the quarter in which a manager rotates. I estimate a specification with worker and quarter fixed effects, binning event time dummy variables at $t=-8$ and $t=4$, and clustering standard errors at the worker and rotation level. Panel A of Figure 3 presents quarterly event study coefficients and demonstrates that manager rotations result in a distinct transitory increase in worker applications. Event time $t=0$ denotes the quarter in which a manager rotates. Application rates increase up to three quarters before the manager rotation takes place, which is when managers start to inform their teams about their departure. However, before $t=-3$, when managers do not know yet about their job rotation, trends in applications are flat. In the quarter of rotation, applications increase by 2.3 percentage points, almost doubling workers’ baseline application rate of 2.7%. As the manager’s replacement settles in, application rates return to baseline levels after one quarter. [^21] Note that applications taking a quarter to return to baseline is consistent with the fact that in a given quarter, only few job openings are available in the internal labor market, implying that not all workers who decide to apply once they find out about the manager rotation will immediately find a suitable job opening.
Since manager rotations appear to have the largest impact on worker applications in the quarter of the rotation, the remainder of the analysis focuses on worker behavior in that quarter. I analyze the effect of manager rotations on worker applications in the same quarter by estimating a linear model for workers’ internal application choices using an OLS regression of the following form:
@@ 5.1 Rotation Design
$\text{Applied}_{it}$ and $\text{Rotation}_{it}$ are indicators that worker $i$ in quarter $t$ applies for an internal job opening and experiences a manager rotation, respectively. $X_{it}$ includes a broad set of worker and position controls. [^22] My baseline set of controls, described in Section 3.2, include female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, and number of direct reports. Appendix Table E11 documents that my results are robust to including different controls. $\delta_{t}$ represents quarter fixed effects. Robust standard errors are two-way clustered at the rotation and worker level.
Manager rotations have large positive impacts on worker applications in the quarter of rotation. Column 1 of Table 5, Panel A presents OLS estimates for the effect of manager rotations on worker applications, based on Equation 3. When a manager rotates, applications increase by 2.3 percentage points, representing a 78% increase compared to the baseline application rate of 2.9%. The application effect is not a result of managers taking their subordinates with them to the new team or of workers replacing managers in their old position: 97% of applications at the firm are for positions outside of the worker’s current team and not to the manager’s new team. In addition, the observed impact does also not depend on the characteristics of the outgoing or incoming managers (Appendix Table B7). If manager rotations reflect the impact of talent hoarding, these findings imply that managerial talent hoarding deters a meaningful share of applicants from exploring job opportunities outside the team. The following subsection presents evidence in favor of this interpretation. [^23] Appendix Section E.3.5 further supports this interpretation by presenting a series of robustness checks that suggest that key alternative explanations, such as correlated shocks, team-specific milestones, worker-manager loyalty, or role-model effects, are unlikely to be main drivers in this setting. Interpreted in this light, this finding accords with workers’ survey responses discussed in Section 2, in which 41% of workers indicate that they are afraid to apply to internal positions due to talent hoarding.
@@ 5.2 Does Talent Hoarding Drive Rotation Effects?
This section provides evidence that talent hoarding is a key mechanism underlying the impacts of manager rotations. The conceptual framework predicts that different types of workers face different degrees of talent hoarding, depending on factors like productivity and manager costs of hoarding. I test these predictions by examining how the effect of manager rotations varies along several of these dimensions. In each case, the analysis includes a continuous interaction between the rotation indicator and a standardized version of the respective proxy for stronger hoarding incentives. This approach allows me to estimate how the effect of manager rotation on application behavior varies with each dimension. The baseline effect of a manager rotation is shown in blue in Panel A of Figure 4, while the estimated change in this effect associated with a one-standard-deviation shift in the characteristic is shown in orange in the respective panels.
Worker Quality.—The conceptual framework in Section 2.2 predicts that managers are more likely to hoard high-productivity workers. Therefore, when managers rotate and talent hoarding temporarily subsides, application effects should be larger for the high-productivity workers who experienced more talent hoarding than their low-productivity peers. To test this prediction, I construct a measure of worker quality as the predicted value from a regression of applicants’ internal hiring probabilities on observable characteristics such as education, tenure, and function. The predicted value of this regression provides an index of worker quality for all workers, weighting worker characteristics by their importance for hiring prospects within the firm. [^24] This definition of worker quality has the advantage that it reflects the values that the firm places on worker qualifications. Robustness exercises in Appendix Section E.3 find similar patterns when using alternative measures such as educational qualifications or past performance.
Panel A shows that the average effect of a manager rotation on application rates is 2.3 percentage points. Panel B of Figure 4 shows that a one-standard-deviation increase in worker quality is associated with a 1.2 percentage point larger increase in applications following a manager rotation ($p$ = 0.000). The positive and statistically significant interaction confirms that higher-quality workers are much more responsive to manager rotations, consistent with the idea that they face stronger hoarding incentives ex ante.
Departure Costs.—Another prediction is that managers hoard talent more when the costs of worker departures are larger. I test this prediction using two measures of departure costs. The first measure is the number of other workers in the team, since a larger team contains more workers that can compensate for a teammate’s departure. Panel C of Figure 4 shows that the effect of manager rotations on applications is significantly larger in smaller teams. The figure plots the interaction of manager rotations with a standardized (negated) measure of team size, so that positive coefficients indicate stronger effects for smaller teams. By including detailed worker and position characteristics, this analysis compares workers with similar qualifications in similar positions, but in teams of different sizes. A one-standard-deviation decrease in team size is associated with a 0.7 percentage point larger increase in applications following a manager rotation ($p$ = 0.008). Measuring departure costs using the average number of days required to fill a worker’s vacated position yields similar results. Panel D of Figure 4 shows that a one-standard-deviation increase in replacement duration is associated with a 0.54 percentage point increase in the application effect following a rotation ($p$ = 0.056). These results support the prediction that workers in roles with higher departure costs are more subject to hoarding behavior.
Talent Hoarding Propensity.—The framework also predicts that managers who are more prone to hoarding talent should have larger effects on application behavior when they rotate. I test this prediction using my measure of managerial hoarding propensity, based on the average deviation between a manager’s private performance ratings and the publicly visible potential ratings they assign to their workers (see Section 4.2 for details).
The impacts of manager rotations are strongly correlated with direct measures of managers’ propensities to hoard talent. Panel E of Figure 4 shows that workers experience significantly larger application effects when a manager with a higher propensity to hoard talent rotates. I interact the manager rotation indicator with a standardized version of the hoarding measure. A one-standard-deviation increase in hoarding propensity is associated with a 0.61 percentage point larger increase in applications ($p$ = 0.033), confirming that workers respond more strongly when a hoarding-prone manager leaves. The framework also predicts that higher-quality workers are hoarded more intensively by hoarding-prone managers. Figure 5 further tests this prediction by focusing on workers whose departing manager falls in the top tercile of the hoarding distribution and then examines how the rotation effect varies with the continuous measure of worker quality. Among workers whose manager had a high propensity to hoard, applications increase by 3.6 percentage points on average following a rotation (baseline effect). A one-standard-deviation increase in worker quality is associated with an additional 1.7 percentage point increase in applications ($p$ = 0.005), demonstrating that the suppressive effect of hoarding-prone managers is concentrated among high-quality workers.
Risk of Retaliation—In survey results (Section 2.1), workers report fear of manager retaliation as a reason they refrain from applying. If workers are more likely to refrain from applying to positions that carry a greater risk of retaliation, applications to these positions should exhibit particularly large increases around rotations. I test this prediction by comparing workers’ applications to job openings that are closer in proximity to their current position (i.e., within versus outside their current division, functional area, and location). Intuitively, the risk of retaliation is larger if managers are more likely to find out about worker applications. Appendix Table B6 shows that manager rotations have much larger and statistically distinguishable effects on applications in close proximity to workers’ current position with respect to all three dimensions, confirming the prediction. [^25] Manager rotations could lead to an opposite prediction on proximity if they were to operate through the visibility channel instead of retaliation. However, while the suppression of visibility is a common form of talent hoarding, manager rotations do not immediately increase visibility since worker evaluations only occur one to two times a year.
External Transitions.—While managers may frequently learn about workers’ applications to internal positions and possibly retaliate or interfere with those applications, features of the institutional setting make this less likely to be the case for applications to positions outside the firm. [^26] In Germany, it is very uncommon that potential employers contact the current firm to inquire about the candidate. Except for a few specific occupations, such as academics, the use of reference letters is also very uncommon, implying that relative to internal transitions, managers have both limited ability to learn about or interfere with external applications. Under talent hoarding, the impact of manager rotations should thus be larger for internal applications than for external applications outside the firm. Because I do not observe external applications, I test this prediction by comparing the effect of manager rotations on internal job transitions within the firm and external job transitions out of the firm. Panel B of Figure 3 shows that manager rotations only increase worker transitions within the firm and not outside of the firm, even though both types of transitions trend identically in quarters prior to the rotation. Columns 1 and 2 of Appendix Table B5 confirm this finding. Manager rotations increase internal transitions by 1.1 percentage points but have negligible effects on external transitions (0.11 percentage points) compared to the identical baseline rate of 0.8% ($p$ of difference = 0.000).